# TopLoc: A Locality Sensitive Hashing Scheme for Trustless Verifiable Inference

### v1

@@ Abstract
Large language models (LLMs) have proven to be very capable, but access to frontier models currently rely on inference providers which introduces trust challenges – how can we be sure that the provider is using the model configuration they claim? We propose TopLoc, a novel method for verifiable inference that addresses this problem. TopLoc leverages a compact locality sensitive hashing mechanism for intermediate activations which can detect unauthorized modifications to models, prompts, or precision with 100% accuracy, achieving no false positives or negatives in our empirical evaluations. Our approach is robust across diverse hardware configurations, GPU types, and algebraic reorderings, which allows for validation speeds up to 100×100$\times$100 × faster than the original inference. By introducing a polynomial encoding scheme, TopLoc minimizes memory overhead of the generated commits by 1000×1000$\times$1000 ×, requiring only 258 bytes of storage per 32 new tokens compared to the 262KB requirement of storing the token embeddings directly for Llama-3.1-8B-Instruct. Our method empowers users to verify LLM inference computations efficiently, fostering greater trust and transparency in open ecosystems and lays a foundation for decentralized, verifiable and trustless compute protocols.
@@ 1 Introduction
Figure 1: An illustration of a process for utilizing verifiable inference for a trustless distributed inference service. Inference Provider A performs batched inferences and generates commits for the computations performed, while Inference Provider B verifies these commits up to 100×100$\times$100 × faster than the time it takes Inference Provider A to generate the responses. Based on the verification outcome, rewards are granted for valid batches and penalties are applied for invalid ones. Further speedup can be obtained for the validator by not checking every batch but instead sampling randomly. Since Provider A does not know which generations will be checked by Provider B, they are incentivized to be honest on all generations to collect the reward and avoid receiving the penalty.
@@
In recent years, large language models (LLMs) have transformed natural language processing, enabling capabilities such as high-quality text generation, advanced dialogue systems, and improved reasoning. (Grattafiori et al., 2024; Team et al., 2024) While many of these models have historically been provided through proprietary, centralized services, an increasing interest in open-source alternatives has emerged. Open-source LLMs, with publicly available weights, promises greater transparency, reproducibility, and community-driven enhancements. However, hosting these open models still requires signficant hardware resources and expertise. Inference providers, entities that run open-source LLMs on their own hardware and expose model outputs via APIs, have risen to meet the demands of users who lack the resources or expertise to operate large-scale inference pipelines themselves.
@@
However, a critical challenge arises in this open ecosystem: trust. Users must trust that an inference provider is faithfully serving the model as advertised, without secret modifications or manipulations. A provider could, for instance, secretly reduce numerical precision to cut costs, fine-tune the model to introduce certain biases, or prepend an undisclosed system prompt to steer the model’s outputs. Without robust verification methods, users can only rely on the provider’s claims, leaving them vulnerable to having their outputs tampered with. (Chen et al., 2023) There is thus a need for verifiable inference, methods of verifying that a certain model and prompt were used during the inference computation.
@@
A standard approach to perform this verification is to use cryptographically verifiable computing methods. (Sun et al., 2024a; Modulus Labs, 2023; Kang et al., 2022; Sun et al., 2023; Ghodsi et al., 2017) However, they are either restrictive in the operations that are supported such that LLMs cannot be used or are currently too computationally expensive to be practically applied to LLM inference computation. One could also introduce fingerprints or watermarks in the model. However, the introduction of fingerprints or watermarks can degrade model performance. Inference providers can also easily detect when a fingerprinting query is being made and thus be able to cheat on the users useful queries.
@@
Another approach is to record the model’s intermediate tensor activations during inference. (Sun et al., 2024b; Zhang et al., 2024) By making these activations available, a third party could independently re-run the model on the same inputs and verify that the intermediate computations match, thus ensuring the authenticity of the inference. However, the direct storage of these intermediate tensors as the commit can be prohibitively expensive at scale. For example, storing the last hidden activations of 100k query consisting of 4096 tokens for Llama-3.1-70B would take 6.7TB.
@@
We propose a method that is able to reduce the storage cost of the commit by more than 1000x while still maintaining the security guarantees of checking intermediate activations. Our method is also robust to GPU non-determinism and algebraic reorderings of the computation, which allows the validation of the commit to be done up to 100x faster than the original inference.
@@
Contribution. In this paper, we present a novel model inference hashing method called TopLoc, which is easy to implement in modern inference engines and efficient enough to be used to generate commits for all model inferences performed by an untrusted compute provider with minimal overhead.
We show that it is able to detect when a different model, prompt or precision was used with 100% accuracy in our experiments with 3 models on thousands of generations. We also show in our experiments that it does not yield any false positives, and is able to validate commits using prefill even if the commits were generated using decode.
@@
We verify that the method is robust to reordering of the computation caused by different GPU types, tensor parallel dimensions as well as different attention kernel implementations.
@@
Furthermore, we propose a method of reducing the memory requirements of storing comparable points by encoding them as a polynomial, allowing our scheme to only require a commit size of 258 bytes for every 32 new tokens generated.
To the best of our knowledge, we are also the first to show empirical results and analysis on the behavior of floating point rounding errors in transformer model inference computations.
Our method has a 100% true positive and true negative rate based on our testing, providing a strong foundation for building trustless compute protocols.
@@ 2 Related Work
Cryptographic Verifiable Computing. Cryptographic Verifiable Computing allows one to verify that a computation was performed correctly on an untrusted computing provider using mathematical proofs. These techniques have been applied to machine learning models and neural networks. (Sun et al., 2024a; Modulus Labs, 2023; Sun et al., 2023; Kang et al., 2022; Ghodsi et al., 2017) However, most of them require the computation be expressed as an arithmetic circuit, limiting the functions that can be used in the model. The translation of the model to an arithmetic circuit also hurts the model quality and makes the proof generation schemes unable to utilize optimized inference engines such as vLLM  [^1] https://github.com/vllm-project/vllm, TensorRT  [^2] https://github.com/NVIDIA/TensorRT and SGLang  [^3] https://github.com/sgl-project/sglang.
[…]
Activation-based validation. SVIP (Sun et al., 2024b) proposes to train a proxy model to map the relationship between the final-layer activations and labels derived from the inference input to produce a fingerprint and then mitigate the ability of an attacker to reverse engineer the proxy model by regularly rotating a secret in the form of an appended vector. However, the security of the scheme requires the retraining of the proxy model with a trusted third party. Moreover, it also requires that the providers are unable to obtain the client secret, which they may obtain by also being a client themselves.
Verisplit (Zhang et al., 2024) proposes the construction of a Merkle tree hash compression of a portion of the activations. However, this compression utilizes a cryptographic hash function, rendering it incompatible with non-deterministic computation caused by GPU kernel scheduling as well as algebraic reorderings of the computations.
@@ 3.1 Inference Modifications
Inference providers often make adjustments to computation methods to optimize for cost, efficiency, or specific commercial goals. While these modifications can make inference more economical and scalable, they may also impact the quality, accuracy, and transparency of the service provided to users. Below, we outline some common methods and the associated incentives behind these changes, as well as their potential implications for users.
Lower precision. Inference providers have an economic incentive to use lower precision formats like fp8 or bf16 as this significantly reduces the compute and memory requirements. While some models and prompts maintain reasonable performance at lower precisions, users may unknowingly experience quality degradation if the provider advertises that they perform the inference in full precision but actually use lower precision.
KVCache compression. The memory requirements for longer conversations with LLMs are dominated by storing intermediate tensors so that they don’t need to be recomputed later. This converts the time complexity of token generation from quadratic to linear, resulting in a large tokens per second increase, but uses more memory. Intermediate tensors can be compressed to enable longer and faster generations with a slightly reduced response quality. (Shi et al., 2024) Providers typically do not advertise their KVCache compression technique, so users may not know what quality to expect.
Altered model weights. Providers have an economic incentive to distill, merge, or prune weights to reduce compute and memory requirements. They may also be incentivized to finetune the models to incorporate bias into the responses. This could lead to users unknowingly using a model that has been significantly altered from the original. Such modifications might lead to degraded model capabilities and biased responses.
Altered system prompt. Providers have an economic incentive to shorten or optimize the user’s prompt to reduce compute costs, as longer prompts require more processing power and memory. Additionally, providers may be incentivized to modify the system prompt to align with their commercial goals, incorporate specific biases, or prioritize certain outcomes. These changes can occur without the user’s knowledge, potentially leading to responses that differ significantly from what the user intended or expected. This lack of transparency can result in reduced trust and suboptimal experiences for users who are unaware of these underlying adjustments.
3.2 Non-Determinism in Model Inference on GPU
Non-determinism in computations performed on GPUs can arise from factors such as operation scheduling and differences in how intermediate results are handled (Monniaux, 2008) (Whitehead and Fit-Florea, 2011).
@@ 3.2 Non-Determinism in Model Inference on GPU
Variations Across GPU Models. Different GPU models often differ in hardware architecture, precision handling, and how they implement mathematical operations. For example, the number of streaming multiprocessors (SMs) on a GPU can affect how parallel operations are scheduled and executed. These differences in computation order or precision handling can result in small numerical discrepancies, particularly in operations like matrix multiplications or summations, which are central to deep learning.
Differences in CUDA Versions. CUDA versions play a significant role in determining the libraries and kernels used during computation. Updates to CUDA often introduce new optimizations or changes in precision handling, which can lead to slight variations in computation results even when the same inputs are provided to the model.
Different Attention Kernel Implementations. Variations in how attention mechanisms are implemented, such as optimizations to improve memory usage or computational speed, can introduce subtle numerical discrepancies. These changes are often made to leverage hardware-specific features but may lead to slightly different outputs.
Different Tensor Parallel (TP) Dimensions. Tensor parallelism divides computations across multiple devices. The partitioning and aggregation of tensors can introduce numerical errors. Operations like summation and normalization, which depend on the order and precision of intermediate calculations, may not be perfectly consistent across different tensor parallel configurations, resulting in variability in the final output.
Prefill vs Decode Modes. Autoregressive models such as LLMs use different computational modes depending on the task. Prefill mode processes input sequences in bulk, while decode mode handles tokens one at a time, typically for tasks like text generation. These modes employ distinct computational paths and optimizations, which can lead to accumulated differences and different outputs.
@@ 3.3 Source of Error in Transformer Models
In bf16 computations of transformer models (Vaswani et al., 2017), most errors arise from the appearance of exact zeros. These zeros are often the result of a catastrophic cancellation (Goldberg, 1991) within the residual stream of the attention layer.
Algorithm 1 Attention block forward
1:Input: Hidden state tensor $h$
2:Output: Updated hidden state tensor
3:
4:$r\leftarrow\texttt{inputLayerNorm}(h)$
[…]
6:$h\leftarrow h+r$triangleright Possibility of catastrophic cancellation
7:$r\leftarrow\texttt{postAttentionLayerNorm}(h)$
8:$r\leftarrow\texttt{MLP}(r)$
9:$h\leftarrow h+r$triangleright Possibility of catastrophic cancellation
10:return $h$
The occurrence of these exact zeros can be non-deterministic. This is because the matrix multiplications involved in the attention and MLP layer are themselves non-deterministic (Golden et al., 2024). Depending on the result of these operations, a value may sometimes lead to an exact zero, and other times it may not.
Interestingly, this behavior reveals a notable property: small values are more susceptible to rounding errors, whereas larger values tend to be represented consistently after algebraic reordering. This insight motivates us to focus on the larger values in a tensor when designing the hash function. By prioritizing these large values, we can reduce the impact of rounding errors and improve the robustness of the hash function.
@@ 4 The TopLoc Algorithm
The TopLoc algorithm encodes and validates the most salient features of a hidden state tensor using a compact, verifiable proof, as detailed in Algorithms 2 and 3. During the inference of the model, we commit to the top-k values in the last hidden state, which can later be verified by the validator recomputing the last hidden states.
An issue with encoding the top-k indices and values directly as the proof is that more storage is actually spent storing the indices. If we allow for tensors of up to 4 billion elements, we would have to store 4 bytes of data for the indices and 2 bytes for the value per point. We can however come up with a scheme that does not store the indices or values at all while still maintaining comparability.
One way is to interpolate a polynomial that passes through the points generated by the top-k indices and values. Given k points, there always exists a unique k - 1 degree polynomial that goes through the points. This k -1 degree polynomial can be represented by k coefficients. We thus only need to store the k coefficients for the proof, each of which consists of 2 bytes.
In order to avoid floating point issues with the polynomial, we interpolate a polynomial congruence in the integer field instead of one in the real numbers. However, this means we need to find a reproduceable unique mapping of the x values into the modulus group as we cannot interpolate a polynomial that yields different y values for the same value of x. We thus need to find a modulus, $m$, such that the function $f(x)=x\mod m$ is injective on the set of indices.
During proof generation (Algorithm 2), the top-$k$ indices and values are extracted from the tensor, and an injective modulus $m$ is computed to uniquely map these indices. The indices and their corresponding values are encoded into a polynomial, which, along with the modulus, forms the proof.
For validation (Algorithm 3), the proof is decoded to retrieve $k$, $m$, and the polynomial. The top-$k$ features are recalculated and compared against the proof by checking for differences in the exponent and mantissa. Validation succeeds if all error thresholds are not exceeded.
@@
4:$(i,v)\leftarrow\texttt{topk}(h,k)$triangleright Find top-$k$ indices and values
@@
6:$i_{m}\leftarrow i\bmod{m}$triangleright Apply injective mapping of indices
7:$P(x)\leftarrow\texttt{InterpolatePolynomialCongruence}(i_{m},v)$triangleright Find polynomial coefficients
8:$p\leftarrow\texttt{encode}(m,P(x))$triangleright Encode the polynomial and injective modulus as the proof
Algorithm 2 TopLoc Proof Generation Algorithm
@@
4:$k,m,P(x)\leftarrow\texttt{decode}(p)$triangleright Decode k, m and the polynomial from the proof
5:$(i,v)\leftarrow\texttt{topk}(h,k)$triangleright Find top-$k$ indices and values
6:$i_{m}\leftarrow i\bmod{m}$triangleright Apply injective mapping of indices
@@
9:     $(e_{p},m_{p})\leftarrow\texttt{extractBits}(P[i_{m}[j]])$triangleright Extract exponent and mantissa bits of $P(i_{m}[j])$
10:     $(e_{v},m_{v})\leftarrow\texttt{extractBits}(v[j])$triangleright Extract exponent and mantissa bits of $v[j]$
@@
18:     return $False$triangleright Validation fails if any threshold is exceeded
@@
20:return $True$triangleright Validation succeeds if all thresholds are satisfied
Algorithm 3 TopLoc Proof Validation Algorithm

### v2

@@ Abstract
Large language models (LLMs) have proven to be very capable, but access to frontier models currently relies on inference providers. This introduces trust challenges: how can we be sure that the provider is using the model configuration they claim? We propose TopLoc, a novel method for verifiable inference that addresses this problem. TopLoc leverages a compact locality-sensitive hashing mechanism for intermediate activations, which can detect unauthorized modifications to models, prompts, or precision with 100% accuracy, achieving no false positives or negatives in our empirical evaluations. Our approach is robust across diverse hardware configurations, GPU types, and algebraic reorderings, which allows for validation speeds significantly faster than the original inference. By introducing a polynomial encoding scheme, TopLoc minimizes the memory overhead of the generated proofs by 1000×1000$\times$1000 ×, requiring only 258 bytes of storage per 32 new tokens, compared to the 262 KB requirement of storing the token embeddings directly for Llama 3.1-8B-Instruct. Our method empowers users to verify LLM inference computations efficiently, fostering greater trust and transparency in open ecosystems and laying a foundation for decentralized, verifiable and trustless AI services.
Locality Sensitive Hashing, Large Language Models, Verifiable Computing
@@ 1 Introduction
In recent years, large language models (LLMs) have transformed natural language processing, enabling capabilities such as high-quality text generation, advanced dialogue systems, and improved reasoning (Grattafiori et al., 2024; Gemma Team et al., 2024). Inference providers, entities that run open-weights LLMs on their own hardware and expose model outputs via APIs, have risen to meet the demands of users who lack the resources or expertise to operate large-scale inference pipelines themselves.
@@
However, a critical challenge arises in this open ecosystem: trust. Users must trust that an inference provider is faithfully serving the model as advertised, without undisclosed modifications. A provider could, for instance, secretly reduce numerical precision to cut costs, fine-tune the model to introduce certain biases, or prepend an undisclosed system prompt to steer the model’s outputs. Without robust verification methods, users can only rely on the provider’s claims, leaving them vulnerable to having their outputs tampered with (Chen et al., 2023). There is thus a need for verifiable inference — methods of verifying that a certain model and prompt were used during the inference computation.
@@
A standard approach to perform this verification is to use cryptographically verifiable computing methods (Sun et al., 2024a; Modulus Labs, 2023; Kang et al., 2022; Sun et al., 2023; Ghodsi et al., 2017). However, they are either restrictive in the operations that are supported such that LLMs cannot be used or are currently too computationally expensive to be practically applied to LLM inference.
@@
Another approach is to record the model’s intermediate activation tensors during inference (Sun et al., 2024b; Zhang et al., 2024). By making these activations available, a third party could independently rerun the model on the same inputs and verify that the intermediate computations match, thus ensuring the authenticity of inference. However, direct storage of these intermediate tensors as the proof can be prohibitively expensive at scale. For example, storing the last hidden activations of 100,000 queries consisting of 4096 tokens for Llama 3.1-70B would take 6.7 terabytes.
@@
In this work, we propose TopLoc, an inference verification method that can reduce the storage cost of the proof by more than 1000x while still maintaining the security guarantees of checking intermediate activations. The method performs locality-sensitive hashing of the intermediate activations, which encodes the top-k values and indices as a polynomial congruence. This polynomial congruence can then be compared against a tensor obtained from recomputation. Our method is also robust to nondeterminism of GPU operations and algebraic reorderings of the computation, which allows the validation of the proof to be done significantly faster than the original inference.
Our contributions are as follows:
@@
  • We present a novel model inference hashing method called TopLoc, which is easy to implement in modern inference engines with minimal overhead.
  • We show that it is capable of detecting when a different model, prompt, or precision was used with 100% accuracy in our experiments.
@@
  • We verify that the method is robust to reordering of the computation caused by different GPU types, tensor parallel dimensions, as well as different attention kernel implementations.
@@
  • We propose a method of reducing the memory requirements of storing comparable points by encoding them as a polynomial, requiring a proof size of only 258 bytes for every 32 new tokens generated.
@@ 2 Related Work
Cryptographic Verifiable Computing. Cryptographic Verifiable Computing allows one to verify that a computation was performed correctly on an untrusted computing provider using mathematical proofs. These techniques have been applied to machine learning models and neural networks (Sun et al., 2024a; Modulus Labs, 2023; Sun et al., 2023; Kang et al., 2022; Ghodsi et al., 2017). However, most of them require the computations to be expressed as an arithmetic circuit, limiting the functions that can be used in the model. The translation of the model to an arithmetic circuit also hurts the model quality and makes the proof generation schemes unable to utilize optimized inference engines such as vLLM111github.com/vllm-project/vllm, TensorRT222github.com/NVIDIA/TensorRT, and SGLang333github.com/sgl-project/sglang.
Moreover, the size of modern LLMs introduces substantial computational overhead for for these methods. zkLLM (Sun et al., 2024a) takes 986 seconds to generate a proof that takes 803 seconds to validate for each inference computation for LLaMa 2-13B. This would mean that a single request that returns 2000 new outputs tokens would require about 23 days to generate the proof for and then 18 days to validate.
Activation-based validation. SVIP (Sun et al., 2024b) proposes training a proxy model to map the relationship between the final layer activations and labels derived from the inference input to produce a fingerprint and then mitigate the ability of an attacker to reverse engineer the proxy model by regularly rotating a secret in the form of an appended vector. However, the security of the scheme requires retraining of the proxy model by a trusted third party. Moreover, it also requires that the providers cannot obtain the client secret, which they may obtain by also being a client themselves.
Verisplit (Zhang et al., 2024) proposes the construction of a Merkle tree hash compression of a portion of the activations. However, this compression utilizes a cryptographic hash function, rendering it incompatible with nondeterministic computation caused by GPU kernel scheduling, as well as algebraic reordering of the computations.
@@ 3.1 Inference Modifications
Inference providers often make adjustments to computation methods to optimize for cost, efficiency, or specific commercial goals. While these modifications can make inference more economical and scalable, they may also impact the quality, and transparency of the service provided to users.
Lower precision. Inference providers might use lower precision formats, such as fp8 (Micikevicius et al., 2022) or bf16 (Kalamkar et al., 2019), which significantly reduces the inference compute and memory requirements.
KV cache compression. Intermediate tensors can be compressed to enable longer and faster generations with a slightly reduced response quality (Shi et al., 2024).
Altered model weights. Providers may distill, merge, or prune weights to reduce compute and memory requirements.
Altered system prompt. Providers could modify the system prompt to align with their commercial goals, incorporate specific biases, or prioritize certain outcomes.
3.2 Nondeterminism in Model Inference on GPU
Nondeterminism in computations performed on GPUs can arise from operation scheduling and differences in how intermediate results are handled (Monniaux, 2008; Whitehead & Fit-Florea, 2011).
@@ 3.2 Nondeterminism in Model Inference on GPU
Furthermore, there are several other causes for variability in computation results when running large language models on GPUs. For example, different GPU models often differ in hardware architecture and precision handling. In addition to that, the CUDA versions determine the libraries and kernels used during computation, and differences in those versions can affect inference. Moreover, subtle variations in how the attention mechanism and other layers are implemented can introduce subtle numerical discrepancies. Lastly, the partitioning and aggregation of tensors when using tensor parallelism can also cause numerical deviations.
@@ 3.3 Source of Error in Transformer Models
In bf16 computations of transformers (Vaswani et al., 2017), most errors arise from the appearance of exact zeros. These zeros are the result of catastrophic cancellations (Goldberg, 1991) within the residual stream of the attention layer. Because the matrix multiplications involved in attention and MLP layers are nondeterministic (Golden et al., 2024), the occurrence of these exact zeros can be nondeterministic.
Interestingly, this behavior reveals a notable property: small values are more susceptible to rounding errors, whereas larger values tend to be represented consistently after algebraic reordering. This insight motivates us to focus on the larger values in a tensor when designing the hash function. By prioritizing large values, we can reduce the impact of rounding errors and improve the robustness of the hash function. We verify this empirically in Section 5.3.
@@ 4 The TopLoc Algorithm
The TopLoc algorithm encodes and validates the most salient features of a hidden state tensor using a compact, verifiable proof, as detailed in Algorithms 1 and 2.
Storing top-$k$ indices and values directly is inefficient, requiring 6 bytes per point: 4 for the index and 2 for the value. However, we can maintain comparability while storing less by interpolating a polynomial that passes through the points generated by the top-k indices and values. Given $k$ points, there always exists a unique $k-1$-degree polynomial that goes through the points which can be represented by $k$ coefficients. We thus only need to store the $k$ coefficients for the proof, each of which consists of 2 bytes.
In order to avoid floating-point issues with the polynomial, we interpolate a polynomial congruence in the integer field instead of in the real numbers. However, this means we need to find a reproducible unique mapping of the x values into the modulus group, as we cannot interpolate a polynomial that yields different $y$ values for the same value of $x$. We thus need to find a modulus, $m$, such that the function $f(x)=x\mod m$ is injective on the set of indices.
During proof generation (Algorithm 1), the top-$k$ indices and values are calculated and an injective modulus $m$ is computed to uniquely map the indices. The indices and their corresponding values are encoded into a polynomial, which, along with the modulus, forms the proof.
Algorithm 1 TopLoc Proof Generation Algorithm
@@
4:  $(i,v)\leftarrow\texttt{topk}(h,k)$Find top-$k$ indices and values}
@@
6:  $i_{m}\leftarrow i\bmod{m}$
7:  $P(x)\leftarrow\texttt{InterpolateModPolynomial}(i_{m},v)$
8:  $p\leftarrow\texttt{encode}(m,P(x))$
For validation (Algorithm 2), the proof is decoded to retrieve $k$, $m$, and the polynomial. The top-$k$ features are recalculated and compared against the proof by checking for differences in the exponent and mantissa. The validation succeeds if the number of exponent mismatches, mean mantissa differences and median mantissa differences are below a set of thresholds.
Algorithm 2 TopLoc Proof Validation Algorithm
@@
4:  $k,m,P(x)\leftarrow\texttt{decode}(p)$
5:  $(i,v)\leftarrow\texttt{topk}(h,k)$Find top-$k$ indices and values}
6:  $i_{m}\leftarrow i\bmod{m}$
@@
9:     $(e_{p},m_{p})\leftarrow\texttt{extractBits}(P[i_{m}[j]])$
10:     $(e_{v},m_{v})\leftarrow\texttt{extractBits}(v[j])$
@@
18:     return $False$
@@
20:  return $True$
Figure 1: When generating the response, we need to perform one prefill for the input tokens and then multiple decodes for each new token generated. When validating, we can pass all the tokens at once and perform just one prefill. Decodes are not efficient on GPUs because they are memory-bound (Agrawal et al., 2024). Notice that the sequential nature of generation causes us to need the decoder blocks multiple times in the generation computation. This increases the total amount of data movement required to perform the computation. The decodes are thus bottlenecked by the time it takes to move data from GPU HBM (High Bandwidth Memory) to SRAM (Shared Memory).
Appendix C details the subroutines and their respective theoretical computational complexities. The corresponding implementations are also available on GitHub444github.com/PrimeIntellect-ai/toploc.